OpenStudy (anonymous):

Find the dimensions and volume of the right circular cylinder of maximum volume inscribed in a sphere with radius 35cm

OpenStudy (anonymous):

Let R = radius of sphere r = radius of cylinder h = height of cylinder V = volume cylinder

OpenStudy (anonymous):

V = πr²h r² = V/(πh) By the Pythagorean Theorem R² = r² + (h/2)² = r² + h²/4 r² = R² - h²/4 r = √(R² - h²/4)

OpenStudy (anonymous):

Set the equations for r² equal. V/(πh) = R² - h²/4 V = πR²h - πh³/4 Take the derivative with respect to h and set it equal to zero to find the critical point(s). dV/dh = πR² - 3πh²/4 = 0 πR² = 3πh²/4 4R²/3 = h² h = 2R/√3

OpenStudy (anonymous):

Take the second derivative to determine the nature of the critical points. d²V/dh² = -2*3πh/4 = -3πh/2 < 0 implies relative maximum Solve for r. r = √(R² - h²/4) r = √{R² - (2R/√3)²/4} = √{R² - R²/3} r = √{(2/3)R²} = R√(2/3) Compute the volume of the cylinder. V = πr²h = π{R√(2/3)}²{2R/√3} = π{(2/3)R²}{2R/√3} V = 4πR³/(3√3) = (4√3)πR³/9

OpenStudy (anonymous):

Thank you :)